# Any Image Restoration with Efficient Automatic Degradation Adaptation

### v1

@@ Abstract
With the emergence of mobile devices, there is a growing demand for an efficient model to restore any degraded image for better perceptual quality. However, existing models often require specific learning modules tailored for each degradation, resulting in complex architectures and high computation costs. Different from previous work, in this paper, we propose a unified manner to achieve joint embedding by leveraging the inherent similarities across various degradations for efficient and comprehensive restoration. Specifically, we first dig into the sub-latent space of each input to analyze the key components and reweight their contributions in a gated manner. The intrinsic awareness is further integrated with contextualized attention in an X-shaped scheme, maximizing local-global intertwining. Extensive comparison on benchmarking all-in-one restoration setting validates our efficiency and effectiveness, i.e., our network sets new SOTA records while reducing model complexity by approximately -82% in trainable parameters and -85% in FLOPs. Our code will be made publicly available at https://github.com/Amazingren/AnyIR.
@@ 1 Introduction
Figure 1: (a) The PSNR comparison under the 3-Degradation All-in-One setting. (b) Average PSNR and parameters comparison (i.e., our AnyIR archives 0.45 dB PSNR improvement with around 82% parameters reduced compared to the SOTA PromptIR [57])
@@
The complexity of current restoration [43, 89, 90, 77, 6] models arises mainly from the diversity of the types of degradation. Consequently, many existing methods are tailored for a specific type of degradation and application with limited generalization to others. A versatile system would require the integration of multiple specialized models, resulting in a complex and heavy framework. Some recent work has shown that a single architecture can address multiple types of degradation, but these approaches lack parameter unification, leading to multiple checkpoints, each for a specific task[89, 6, 77, 10]. Although this reduces system complexity, it compromises efficiency due to the increased number of checkpoint sets.
Recent studies have focused on achieving both architectural and parameter unification [33, 57, 87, 93, 47]. Diffusion-based methods have demonstrated improved image generation capabilities [62, 29, 99]. To improve the controllability of the network, other approaches treat each modality as an independent visual prompt, which guides the network toward the target application [74, 57, 42]. Another group of work uses text as an intermediate representation to unify all degradations, using textual instructions to prompt the network [49, 11]. Despite promising results, these networks still require significant computational resources, with a huge number of parameters and low inference speed.
@@
In summary, our contributions are threefold:
  • We propose AnyIR, a single model capable of efficiently handling any type of degradation. Compared to the SOTA all-in-one counterparts, our model reduces the computational cost by $85.6\%$ while delivering superior overall performance.
@@ 2 Related Work
Image Restoration (IR), a longstanding and ill-posed inverse problem, aims to reconstruct high-quality images from their degraded counterparts. Due to its practical value, IR has been applied to various real-life applications [65, 2, 41, 88]. Initially, IR was addressed through model-based solutions, involving the search for solutions to specific formulations. However, learning-based approaches have gained much attention with the significant advancements in deep neural networks. Numerous approaches have been developed, including regression-based [44, 32, 43, 7, 39, 97] and generative model-based pipelines [22, 76, 50, 86, 99, 45] that built on convolutional [15, 95, 94, 75], MLP [72], state space mode [24, 100, 23, 14], or vision transformers-based (ViTs) architectures [43, 39, 89, 60, 17]. Though promising performance has been achieved by many previous state-of-the-art methods, the mainstream IR solutions still focus on addressing single degradation tasks such as denoising [95, 98], dehazing [64, 80], deraining [28, 61], deblurring [31, 62], and so on. In this paper, our method is built based on the regression-based pipeline with ViTs architecture.
Figure 2: (a) The overall framework of the proposed AnyIR is in a U-shaped structure which mainly consists of a convolutional patch embedding, a U-shape encoder-decoder main body, and an extra refined block. (b) The structure of each degradation adaptation block (DAB)
@@
Training a task-specific model to tackle a single type of degradation is effective, but practical implementation is challenging due to the need for separate models for each degradation type. In reality, images often suffer from multiple degradations and artifacts, making it difficult to address them individually. Task-specific solutions require significant computing and storage resources, substantially increasing their environmental impact. To address this, the emerging All-in-One image restoration field uses a single blind restoration model to handle multiple degradation types simultaneously. Specifically, AirNet [34] accomplishes blind All-in-One image restoration by leveraging contrastive learning to derive degradation representations from corrupted images, which are then used to reconstruct the clean image. Following this, IDR [93] addresses All-in-One image restoration by decomposing image degradations into their fundamental physical principles and employing a two-stage meta-learning approach. More recently, the prompt-based paradigm [57, 74, 42] has been introduced, incorporating an additional visual prompt learning module. This approach has been validated as effective in guiding a single trained model to better restore images with multiple types of degradation by leveraging the learned discriminative capabilities of the visual prompts. Built upon the prompt-based paradigm, there also some recent works propose to modeling the visual prompt either from the frequency perspective [12] or by posting more complex architecture designs with extra dataset [18, 87]. However, using visual prompt modules often leads to longer training times and reduced efficiency. Instead, our work enhances the model’s ability to capture representative information without the burden of heavy and complex prompt modules.
3 Proposed AnyIR
In this section, we present our efficient All-in-One restoration method, termed AnyIR. The overall framework is illustrated in Fig. 2. At a macro level, AnyIR employs a U-shaped network architecture [66] with four levels. Each level incorporates $N_{i},i\in[1,2,3,4]$ instances of our novel Degradation Adaptation Blocks (DAB) 3.1, and each DAB is built upon the proposed gated degradation adaption (GatedDA) module (Sec. 3.2). Initially, a convolutional layer first extracts shallow features from the degraded input, creating a patch embedding of $H\times W\times C$. Subsequently, each U-Net level doubles the embedding dimension and halves the spatial resolution. Skip connections transfer information from the encoder stage to the decoder. At the decoder, we merge these features with the signal from the previous decoding stage using a linear projection. Lastly, a global residual connection links the input image to the output, preserving high-frequency details and producing the restored image.
3.1 Degradation Adaptation Module
@@
The detailed structure of the proposed DAM is shown in Fig. 2 (b). Given the feature with $F_{in}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, width, and channel numbers, respectively. We first split it along the channel dimension with a skip selection strategy as follows:
$F_{in}^{att}=F_{in}^{2i-1},F_{in}^{gate}=F_{in}^{2i},$ (1)
$F_{in}^{att}$, $F_{in}^{gate}$ $\in\mathbb{R}^{H\times W\times\nicefrac{{C}}{{2}}}$ denote the divided features with $i\in[1,2,3,\cdots,\nicefrac{{C}}{{2}}]$. This simple skip-split operation offers two significant advantages. (i) The model complexity is greatly reduced as the channel dimension is halved, which substantially decreases the number of model parameters, particularly within the self-attention mechanism of ViTs [17]. (ii) Unlike the commonly adopted half-split method, the proposed skip-split ensures that each divided half-feature originates from uniform channel-wise downsampling, thus preventing severe information loss in each split.
Then multi-depth convolution head attention [89, 57] is applied to $F_{in}^{att}$ for modeling the global degradation dependence and output $F_{out}^{att}$, i.e. $F_{out}^{att}=\sum_{i}\texttt{Softmax}(\nicefrac{{\mathbf{Q}\mathbf{K}}}{{\sqrt{d}}})_{i}\mathbf{V}_{i,j}$, where $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ are query, key and value matrices. $d$ denotes the dimension of $\mathbf{Q}$ and $\mathbf{K}$. Meanwhile, the proposed GatedDA (Sec. 3.2) operation is applied to $F_{in}^{gate}$ and outputs $F_{out}^{gate}$. This design leverages the inherent properties of the attention mechanism to capture crucial structural cues for restoring the original image quality. Concurrently, we utilize the strengths of convolutional operators to extract local dependencies, which are vital in low-level vision applications [43, 9, 8].
To ensure that both global and specialized local features benefit each other, we propose a cross-feature update as follows:
$F_{out}^{ag}$ $=F_{out}^{att}+\operatorname{Sigmod}(F_{out}^{gate}),$ (2)
$F_{out}^{ga}$ $=F_{out}^{gate}+\operatorname{Sigmod}(F_{out}^{att}),$
here, $F_{out}^{ag}$ and $F_{out}^{ga}$ denote the mutually beneficial features. For cross-feature filtering, we first apply Sigmoid activation to both intermediate features. Each activated signal is then added to the corresponding opposite output. Finally, we use a linear projection followed by a feed-forward network (FFN) and a residual connection to generate the output $F_{out}$.
@@ 3.2 Gated Degradation Adaption
To better capture the local degradation-aware details, we propose to utilize the natural selective properties of the gated convolutional operation. Specifically, given an input feature $F_{in}\in\mathbb{R}^{H\times W\times C}$, we first normalize it with batch normalize operation. Then we use the convolutional operation with 1 $\times$ 1 kernel size, to linearly project the channel dimension $C$ to a hidden dimension $hidden$. And we set the $hidden=2\times rC$. $r_{expan}$ denotes the expansion ratio. We formalize this process as follows:
$F^{\prime}=\mathbf{W}_{1}(\text{Norm}(F_{in})),$ (3)
here $\mathbf{W}_{1}$ is the learnable parameters with expansion ratio $r_{expan}$. $F^{\prime}$ denotes the channel-wise linear convolutional projection layer.
Then $F^{\prime}$ is split into three parts, i.e., $\alpha$ (shifted), $\beta$ (ego), and $\gamma$ (scaled) along the hidden channel dimension with the channel dimension $r_{dconv}\times C$, ($\nicefrac{{hidden}}{{2}}-r_{dconv})\times C$, and $\nicefrac{{hidden}}{{2}}$. Then $\alpha$ is set to the depth-wise convolutional operation for capturing the spatial-wise details, $\beta$ is kept without any operation to maintain the original of $F^{\prime}$, and $\gamma$ is passed through a GELU [25] based activation function for constructing the gated selection mechanism. This process is formalized as:
$F^{\prime\prime}=\text{Cat}(\mathbf{W}_{2}\alpha,\beta)\times\text{GELU}(\gamma),$ (4)
where Cat(.), $\mathbf{W}_{2}$, and GELU denote the channel-wise concatenation, the learnable parameters of the depth-wise convolutional filter, and the non-linear GELU [25] activation function, respectively. GELU here introduces nonlinearity, enabling networks to better fit complex function relationships. And the multiplication here leads to a modulation of the contacted feature which helps enhance the model’s expressive power and performance.
Then, we project $F^{\prime\prime}$ back to its original channel dimension via another channel-wise linear convolutional projection layer with one skip connection linked to the input feature to prevent information loss. Finally, to well fuse the original features $F_{in}$ and the degradation adapted feature, another convolutional filter with kernel size 1 is used. We formulate this process as follows:
$F_{out}=\mathbf{W}_{4}(\mathbf{W}_{3}F^{\prime\prime}+F_{in}),$ (5)
where $\mathbf{W}_{3}$ and $\mathbf{W}_{4}$ are the learnable parameters of the convolutional filters.
Figure 3: The structure of our GatedDA. $\oplus$, circled-c, circled-g, and $\otimes$ denote the element-wise addition, channel-wise concatenation, GELU [25] activation, and element-wise multiplication, respectively.
@@
AnyIR (Ours) 6M 31.38 .979 37.90 .981 33.95 .933 31.29 .889 28.03 .797 32.51 .916

### v2

# Restore Anything Model via Efficient Degradation Adaptation

@@ Abstract
With the proliferation of mobile devices, the need for an efficient model to restore any degraded image has become increasingly significant and impactful. Traditional approaches typically involve training dedicated models for each specific degradation, resulting in inefficiency and redundancy. More recent solutions either introduce additional modules to learn visual prompts—significantly increasing model size—or incorporate cross-modal transfer from large language models trained on vast datasets, adding complexity to the system architecture. In contrast, our approach, termed RAM, takes a unified path that leverages inherent similarities across various degradations to enable both efficient and comprehensive restoration through a joint embedding mechanism—without scaling up the model or relying on large multimodal models. Specifically, we examine the sub-latent space of each input, identifying key components and reweighting them in a gated manner. This intrinsic degradation awareness is further combined with contextualized attention in an X-shaped framework, enhancing local-global interactions. Extensive benchmarking in an all-in-one restoration setting confirms RAM’s SOTA performance, reducing model complexity by approximately 82% in trainable parameters and 85% in FLOPs. Our code and models will be publicly available.
@@ 1 Introduction
Figure 1: Average PSNR and parameters comparison (i.e., our RAM archives 0.45 dB PSNR improvement with around 82% parameters reduced compared to the SOTA PromptIR [53])
@@
The complexity of existing restoration models [42, 86, 84, 73, 5] arises mainly from the diversity of the types of degradation. Consequently, many existing methods are tailored for a specific type of degradation and application with limited generalization to others. A versatile system would require the integration of multiple specialized models, resulting in a complex and heavy framework. Some recent work has shown that a single architecture can address multiple types of degradation, but these approaches lack parameter unification, leading to multiple checkpoints, each for a specific task[86, 5, 73, 8]. Although this reduces system complexity, it compromises efficiency due to the increased number of checkpoint sets.
Recent studies have focused on achieving both architectural and parameter unification [34, 53, 82, 88, 46]. Diffusion-based methods have demonstrated improved image generation capabilities [58, 30, 95]. To improve the controllability of the network, other approaches treat each modality as an independent visual prompt, which guides the network toward the target application [70, 53, 41]. Another group of works uses text as an intermediate representation to unify all degradations, using textual instructions to prompt the network [49, 12]. Despite promising results, these networks still require significant computational resources, with a huge number of parameters and low inference speed.
@@
The contributions of this work are:
  • We propose RAM , a single model capable of efficiently handling any type of degradation. Compared to the SOTA all-in-one counterparts, our model reduces the computational cost by $85.6\%$ while delivering superior overall performance.
@@ 2 Related Work
Image Restoration (IR). Image restoration aims to solve a highly ill-posed problem by reconstructing high-quality images from their degraded counterparts. Due to its importance, IR has been applied to various applications [61, 1, 40, 83]. Initially, IR was addressed through model-based solutions involving the search for solutions to specific formulations. However, learning-based approaches have gained much attention with the significant advancements in deep neural networks. Numerous approaches have been developed, including regression-based [43, 32, 42, 7, 39, 92] and generative model-based pipelines [24, 72, 48, 81, 95, 44] that built on convolutional [16, 90, 89, 71], MLP [68], state space mode [26, 96, 25, 15], or vision transformers-based (ViTs) architectures [42, 39, 86, 55, 18]. Although state-of-the-art methods have achieved promising performance, the mainstream IR solutions still focus on addressing single degradation tasks such as denoising [90, 94], dehazing [60, 76], deraining [29, 57], deblurring [31, 58], and so on. In this work, the proposed method is built based on the regression-based pipeline with ViT architecture.
Figure 2: (a) Overall framework of the proposed RAM : a U-shaped structure which mainly consists of a convolutional patch embedding, a U-shape encoder-decoder main body, and an extra refined block. (b) Structure of each degradation adaptation block (DAB)
@@
Training a task-specific model to tackle a single type of degradation is effective, but practical implementation is challenging due to the need for separate models for each degradation type. In reality, images often suffer from multiple degradations and artifacts, making it difficult to address them individually. Task-specific solutions require significant computing and storage resources, substantially increasing their environmental impact. To address this, the emerging All-in-One image restoration field uses a single-blind restoration model to handle multiple degradation types simultaneously. Specifically, AirNet [35] accomplishes blind All-in-One image restoration by leveraging contrastive learning to derive degradation representations from corrupted images, which are then used to reconstruct the clean image. Following this, IDR [88] addresses All-in-One image restoration by decomposing image degradations into their fundamental physical principles and employing a two-stage meta-learning approach. More recently, the prompt-based paradigm [53, 70, 41] has been introduced, incorporating an additional visual prompt learning module. This approach has been validated as effective in guiding a single trained model to better restore images with multiple types of degradation by leveraging the learned discriminative capabilities of the visual prompts. Built upon the prompt-based paradigm, there also some recent works propose modeling the visual prompt either from the frequency perspective [13] or by posting more complex architecture designs with extra dataset [20, 82]. However, using visual prompt modules often leads to longer training times and reduced efficiency. Instead, our work enhances the model’s ability to capture representative information without the burden of heavy and complex prompt modules.
3 Proposed RAM
Overview.
In this section, we present our efficient All-in-One restoration method, termed RAM . The overall framework is illustrated in Fig. 2. At a macro level, RAM employs a U-shaped network architecture [62] with four levels. Each level incorporates $N_{i},i\in[1,2,3,4]$ instances of our novel Degradation Adaptation Blocks (DAB) 3.1, and each DAB is built upon the proposed gated degradation adaption (GatedDA) module (Sec. 3.2). Initially, a convolutional layer first extracts shallow features from the degraded input, creating a patch embedding of $H\times W\times C$. Subsequently, each U-Net level doubles the embedding dimension and halves the spatial resolution. Skip connections transfer information from the encoder stage to the decoder. At the decoder, we merge these features with the signal from the previous decoding stage using a linear projection. Lastly, a global residual connection links the input image to the output, preserving high-frequency details and producing the restored image.
3.1 Degradation Adaptation Block
@@
The detailed structure of the proposed DAB is illustrated in Fig. 2 (b). Given an input feature tensor $F_{in}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, width, and channel dimensions, respectively, we employ a selective channel-wise partitioning strategy defined as follows:
$F_{in}^{att}$ =$=\{F_{in}^{(2k-1)}\mid k\in\mathbb{Z}^{+},k\leq\frac{C}{2}\},$ (1)
$F_{in}^{gate}$ =$=\{F_{in}^{(2k)}\mid k\in\mathbb{Z}^{+},k\leq\frac{C}{2}\},$
where $F_{in}^{att}$ and $F_{in}^{gate}\in\mathbb{R}^{H\times W\times\nicefrac{{C}}{{2}}}$ denote the sub-features derived from interleaved sampling along the channel dimension. This setup provides two distinct pathways: $F_{in}^{att}$, focusing on attention modulation, and $F_{in}^{gate}$, specializing in gated transformation. This structured skip-split offers two notable benefits, i.e., (i) Significant Complexity Reduction: By reducing the effective channel dimension in each pathway, the overall model complexity and parameter count are notably minimized, which is particularly beneficial for the self-attention layers in Vision Transformers (ViTs) [18]. (ii) Improved Information Retention: Unlike conventional half-split methods, the skip-split preserves feature diversity by uniformly sampling the original channels, ensuring each subset retains comprehensive information from $F_{in}$ without loss of essential channel-wise details.
Figure 3: Structure of our GatedDA. $\oplus$, circled-c, circled-g, and $\otimes$ denote the element-wise addition, channel-wise concatenation, GELU [27] activation, and element-wise multiplication, respectively.
To capture the complex dependencies inherent to global degradation, we employ a multi-depth convolution head attention mechanism [86, 53] on the feature subset $F_{in}^{att}$, resulting in $F_{out}^{att}$. This attention approach is particularly effective for image restoration, where degradation is often non-uniform and shaped by various intricate factors. Specifically, $F_{in}^{att}$ is transformed into Query ($Q$), Key ($K$), and Value ($V$) matrices, defined as: $Q=F_{in}^{att}\mathbf{W}_{qry},\quad K=F_{in}^{att}\mathbf{W}_{key},\quad V=F_{in}^{att}\mathbf{W}_{val},$ where $\mathbf{W}_{qry}$, $\mathbf{W}_{key}$, and $\mathbf{W}_{val}$ are learnable weights. The output $F_{out}^{att}$ is then calculated as:
$F_{out}^{att}=\sum_{i}\texttt{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)_{i}V_{i,j},$ (2)
where $\sqrt{d}$ is a scaling factor to normalize attention scores, stabilizing gradients and aiding convergence. This attention mechanism excels at modeling long-range dependencies across the feature space, a vital capability for image restoration where degradation—such as noise, blur, and artifacts—exhibits spatial correlations yet varies across the image [42, 86, 39, 6]. By leveraging this global modeling, the framework dynamically identifies and emphasizes degraded regions, preserving fine details essential for high-quality restoration.
Meanwhile, the proposed GatedDA operation (see Sec. 3.2) is applied to $F_{in}^{gate}$, producing $F_{out}^{gate}$. This approach combines the strengths of attention and convolution: the attention module captures global structural cues essential for overall image context, while the convolutional pathway captures local dependencies crucial for refining details in low-level vision tasks [42, 10, 9]. To enable effective interaction between these global and local features, we introduce a cross-feature update:
$F_{out}^{ag}$ $=F_{out}^{att}+\sigma\left(F_{out}^{gate}\right),$ (3)
$F_{out}^{ga}$ $=F_{out}^{gate}+\sigma\left(F_{out}^{att}\right),$
where $\sigma(\cdot)$ is the Sigmoid activation, enhancing each feature map with complementary information. Through this update, $F_{out}^{att}$ gains local detail sensitivity, while $F_{out}^{gate}$ benefits from structural context. The fused representation $F_{fuse}$ is then computed by concatenating $F_{out}^{att}$ and $F_{out}^{gate}$, followed by a convolution with learnable weights $\mathbf{W}_{fuse}$ and a residual connection to the input:
$F_{fuse}=\operatorname{Conv}_{\mathbf{W}_{fuse}}(\operatorname{concat}(F_{out}^{att},F_{out}^{gate}))+F_{in}.$ (4)
This fusion step integrates global and local information, enhancing feature richness for the final restoration. Next, layer normalization, a feed-forward network (FFN), and a residual connection are applied, yielding $F_{out}$:
$F_{out}=\operatorname{FFN}(\operatorname{Norm}(F_{fuse}))+F_{fuse}.$ (5)
This sequence stabilizes feature distributions and boosts expressive capacity, effectively combining global and local cues for high-fidelity, detailed image restoration.
@@ 3.2 Gated Degradation Adaption
Algorithm 1 Gated Degradation Adaption
1:Input $F_{in}^{\text{gate}}\in\mathbb{R}^{C\times H\times W}$, initial temperature $\tau$
2:Output $F_{out}^{\text{gate}}\in\mathbb{R}^{C\times H\times W}$
3:$\hat{F}\leftarrow\text{Norm}(F_{in}^{\text{gate}})$ // Normalize
4:$\mu,\sigma\leftarrow\frac{1}{HW}\sum\hat{F},\sqrt{\frac{1}{HW}\sum(\hat{F}-\mu)^{2}}$ // Mean, std
5:$\alpha\leftarrow\sigma(\mu+\sigma)$ // Temp adjust
6:$\tau_{\text{adj}}\leftarrow\tau\cdot\alpha$
7:$F^{\prime}\leftarrow\hat{F}\mathbf{W}_{\text{exp}}$ // Channel expand
8:$\gamma,\beta,\alpha\leftarrow\text{split}(F^{\prime})$ // Split $\gamma$, $\beta$, $\alpha$
9:$\alpha^{\prime}\leftarrow(\alpha\mathbf{W}_{\text{depth}})\cdot\tau_{\text{adj}}$ // Depthwise conv
10:$F_{\text{gated}}\leftarrow\sigma(\gamma)\cdot\text{concat}(\beta,\alpha^{\prime})\mathbf{W}_{\text{gate}}$ // Gate combine
11:$F_{out}^{\text{gate}}\leftarrow(F_{\text{gated}}+F_{in}^{\text{gate}})\mathbf{W}_{\text{proj}}$ // Residual, proj
12:return $F_{out}^{\text{gate}}$
@@
RAM (Ours) 6M 31.38 .979 37.90 .981 33.95 .933 31.30 .889 28.04 .797 32.51 .916
Methods with the assistance of vision language, multi-task learning, natural language prompts, and multi-modal control
DA-CLIP [49] 125M 29.46 .963 36.28 .968 30.02 .821 24.86 .585 22.29 .476 - -
ArtPromptIR [75] 36M 30.83 .979 37.94 .982 34.06 .934 31.42 .891 28.14 .801 32.49 .917
InstructIR-3D [11] 16M 30.22 .959 37.98 .978 34.15 .933 31.52 .890 28.30 .804 32.43 .913
UniProcessor [19] - 31.66 .979 38.17 .982 34.08 .935 31.42 .891 28.17 .803 32.70 .918
To capture local degradation-aware details, we leverage the selective properties of gated convolution. Given an input feature map $F_{in}^{\text{gate}}\in\mathbb{R}^{C\times H\times W}$, a layer normalization is first applied to stabilize feature distributions, followed by a 1$\times$1 convolution that expands the channel dimension to $\text{hidden}=r_{\text{expan}}\cdot C$, where $r_{\text{expan}}$ is the expansion ratio, as shown in Alg. 1. To adaptively respond to varying intensities, we introduce a temperature adjustment mechanism. Based on the input’s mean and standard deviation, the initial temperature $\tau$ is modulated as $\tau_{\text{adj}}=\tau\cdot\alpha$, where $\alpha$ serves as a dynamic scaling factor (see steps 2–4 of Algorithm 1). This temperature adjustment enhances the model’s sensitivity to subtle degradation patterns, promoting detailed feature capture. The expanded feature $F^{\prime}$ is then split into three parts: $\gamma$, $\beta$, and $\alpha$ (as shown in step 6). Here, $\alpha$ undergoes depthwise convolution with $\tau_{\text{adj}}$ to capture spatial details, $\beta$ retains the original information, and $\gamma$ is activated with GELU [27] to enable a non-linear gated selection mechanism. These components are recombined and modulated, selectively emphasizing critical features (step 7). Finally, $F^{\prime\prime}$ is projected back to the original channel dimension and combined with $F_{in}^{\text{gate}}$ through a skip connection to prevent information loss. A final 1$\times$1 convolution fuses the degradation-adapted features with the input, yielding the output $F_{out}^{\text{gate}}$ (steps 8–9 in Algorithm 1). In summary, the GatedDA block dynamically adjusts temperature and gating mechanisms to adaptively capture degradation-sensitive features, balancing global structure with fine detail. Further analysis is available in the Supplementary Materials (i.e., Supp. Mat.).